# Kernel Ridge Regression with Predicted Feature Inputs and Applications to Factor-Based Nonparametric Regression

### v1

@@ 1.2.1 Non-asymptotic excess risk bounds of KRR with predicted feature inputs
1.2.2 A new framework for analyzing KRR with predicted feature inputs
Along with our new rate results comes a novel approach to analyzing the KRR with predicted inputs. The framework we develop offers two key contributions: (1) it improves existing analyses of KRR in the classical setting by accommodating arbitrary model misspecification and allowing both an unbounded response and an unbounded RKHS; (2) it is essential for handling predicted feature inputs, enabling us to derive risk bounds without imposing any assumptions on the proximity of $\widehat{g}(X)$ to $Z$.
To accomplish the first point, we adopt a hybrid approach that combines the integral operator method with the local Rademacher complexity based empirical process theory from Bartlett et al. (2005), along with a careful reduction argument. This is necessary because the integral operator approach cannot easily handle model misspecification, while the empirical process theory requires both a bounded response and a bounded RKHS. Specifically, as seen in the decomposition (8), the key is to bound from above $\mathbb{E}[\ell_{\widehat{f}\circ\widehat{g}}(Y,X)]$. Since $\sup_{f\in\mathcal{H}_{K}}\|f\|_{K}$ could be unbounded, we first apply a reduction argument, showing that bounding $\mathbb{E}[\ell_{\widehat{f}\circ\widehat{g}}(Y,X)]$ can be reduced to controlling the empirical process
@@ 1.2.2 A new framework for analyzing KRR with predicted feature inputs
where $\uline{\mathcal{F}_{b}}$ is a bounded subset of $\mathcal{H}_{K}$, given by (37) of Section 3.4. We use $\mathbb{E}_{n}$ in this paper to denote the expectation with respect to the empirical measure of $n$ i.i.d. samples. To further address the unboundedness of $Y$, we need to bound the cross-term $\mathbb{E}_{n}[\epsilon(f\circ\widehat{g})(X)-\epsilon(f_{\mathcal{H}}\circ\widehat{g})(X)]$ uniformly over $f\in\uline{\mathcal{F}_{b}}$. To accomplish this, we employ both the empirical process theory based on local Rademacher complexity from Bartlett et al. (2005) and the integral operator approach, as detailed in Section B.2.3; see also Lemma 7. Our analysis only requires moment conditions on the regression error $\epsilon$. Finally, the local Rademacher complexity technique from Bartlett et al. (2005) is applied once again to control the remaining bounded components in the empirical process (10). Due to our new hybrid approach, the analysis can be easily applied to learning algorithms in (1) in which other convex loss functions are used in place of the squared loss, as demonstrated in Section 5.
@@
Another fundamental difficulty in our analysis arises from handling predicted feature inputs. As noted in the end of Section 1.1, the primary challenge lies in establishing a precise connection between the local Rademacher complexity and the kernel complexity function $R(\delta)$ in the presence of errors in predicting the features. The existing analysis in Mendelson (2002) only establishes a connection to $R_{x}(\delta)$, which depends on the probability measure $\rho_{x}$ of $\widehat{g}(X)$. And it is hopeless to relate $R_{x}(\delta)$ with $R(\delta)$ without imposing restrictive assumptions on the closeness of $\rho_{x}$ and $\rho$. To circumvent this issue, rather than working with the population-level local Rademacher complexity and $R_{x}(\delta)$—both of which rely on $\rho_{x}$—we instead focus on their empirical counterparts. The first step is to relate the local Rademacher complexity to its empirical counterpart where we borrow existing concentration inequalities of local Rademacher complexities in Boucheron et al. (2000, 2003); Bartlett et al. (2005). By adapting the argument in Bartlett et al. (2005, Lemma 6.6), we further link the empirical local Rademacher complexity to $\widehat{R}_{x}(\delta)$, the empirical counterpart of $R_{x}(\delta)$. The next step is to connect $\widehat{R}_{x}(\delta)$ to $\widehat{R}(\delta)$, with the latter being the empirical counterpart of $R(\delta)$. This is the most challenging step, as $\widehat{R}_{x}(\delta)$ and $\widehat{R}(\delta)$ depend, respectively, on the empirical integral operators associated with $\widehat{g}(X_{i})$ and $Z_{i}$ for $i\in[n]$. Our proof reveals that this requires to control the spectral difference between two $n\times n$ kernel matrices with entries equal to $K(\widehat{g}(X_{i}),\widehat{g}(X_{j}))$ and $K(Z_{i},Z_{j})$ for $i,j\in[n]$. As accomplished in Lemmas 11 and 12 of Section B.3.3, this is highly non-trivial as we do not impose any requirements on the proximity of $\widehat{g}(X_{i})$ to $Z_{i}$. Finally, we establish sharp concentration inequalities between $\widehat{R}(\delta)$ and its population-level counterpart, $R(\delta)$, to close the loop of relating the local Rademacher complexity to $R(\delta)$. Completing the loop requires a series of technical lemmas, collected in Lemmas 9, 10, 11, 12, and 13 of Sections B.3.1, B.3.2, B.3.3 and B.3.4. Developing this proof strategy is part of our contribution, and we provide a more detailed discussion in Section 3.4.3. For illustration, Figure 1 below depicts how we relate the local Rademacher complexity to the kernel complexity $R(\delta)$, represented by the solid arrow.
@@
1.2.3 Application to factor-based non-parametric regression models
Our third contribution is to provide a complete analysis of using KRR with feature inputs predicted by PCA under the factor-based nonparametric regression models (3) and (4). As an application of our general theory, together with the existing literature on high-dimensional factor models (for instance, Bai (2003); Fan et al. (2013)), we state in Corollary 5 of Section 4.2 an explicit upper bound of the excess risk of the KRR with inputs predicted by PCA. Furthermore, in Theorem 3 of Section 4.2, we establish a matching minimax lower bound of the excess risk under models (3) and (4), thereby demonstrating the minimax optimality of the KRR using the leading PCs.
@@ 1.2.3 Application to factor-based non-parametric regression models
This paper is organized as follows: we introduce the learning algorithm of KRR with predicted inputs in Section 2. Theoretical results of analyzing KRR with predicted inputs are stated in Section 3. The risk decomposition is given in Section 3.1 while the risk bounds along with the main assumptions are stated in Section 3.2. Risk bounds for specific kernels are given in Section 3.3 while in Section 3.4 we explain the proof sketch and highlight the main technical difficulties. In Section 4 we apply our general result to factor-based nonparametric regression models. In Section 5 we extend our analysis and derive risk bounds for general convex loss functions. Simulation studies are stated in Appendix A and all proofs are deferred to the Appendix.
Notation. For any $a,b\in\mathbb{R}$, we write $a\wedge b=\min\{a,b\}$ and $a\vee b=\max\{a,b\}$. For any integer $d$, we let $[d]=\{1,\ldots,d\}$. For any symmetric, semi-positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we use $\lambda_{1}(Q)\geq\lambda_{1}(Q)\geq\ldots\geq\lambda_{d}(Q)$ to denote its eigenvalues. For any matrix $A$, we use $\|A\|_{\rm op}$ and $\|A\|_{\text{F}}=(\sum A^{2}_{ij})^{1/2}$ to denote its operator norm and Frobenius norm, respectively. We use $\mathbf{I}_{d}$ to denote the $d\times d$ identity matrix. For $d_{1}\geq d_{2}$, we write $\mathbb{O}_{d_{1}\times d_{2}}$ for matrices with orthonormal columns. The inner product and the endowed norm in the Euclidean space are denoted as $\langle\cdot,\cdot\rangle_{2}$ and $\|\cdot\|_{2}$, respectively. For any function $f:\mathcal{Z}\to\mathbb{R}$, $\|f\|_{\infty}=\sup_{z\in\mathcal{Z}}|f(z)|$. For any two sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C$ such that $a_{n}\leq Cb_{n}$ for all $n$. We write $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. In this paper, we use $c,c^{\prime},C$ and $C^{\prime}$ to denote constants whose values may vary line by line unless otherwise stated.
@@ 2 Kernel ridge regression with predicted feature inputs
In this section, we state the learning algorithm for Kernel Ridge Regression (KRR) with predicted feature inputs. Recall that $\mathcal{D}=\{(Y_{i},X_{i})\}_{i=1}^{n}\in(\mathbb{R},\mathcal{X})^{n}$ contains $n$ training data. Let $\widehat{g}:\mathcal{X}\to\mathcal{Z}$ be any measurable function that is used to predict the features $Z_{i}\in\mathcal{Z}$. For instance, $\widehat{g}$ could be obtained from one of the feature extraction or representation learning approaches mentioned in the Introduction. The predicted feature inputs are $\widehat{g}(X_{i})$ for all $i\in[n]$.
Let $K:{\mathcal{Z}\times\mathcal{Z}}\to\mathbb{R}$ be some pre-specified kernel function and denote by $\mathcal{H}_{K}$ the Reproducing Kernel Hilbert Space (RKHS) induced by $K$. Write the equipped inner product in $\mathcal{H}_{K}$ as $\langle\cdot,\cdot\rangle_{K}$ and the endowed norm as $\|\cdot\|_{K}$. For any $v\in\mathcal{Z}$, by writing $K_{v}:=K(v,\cdot)\in\mathcal{H}_{K}$, the reproducing property asserts that
$\langle f,K_{v}\rangle_{K}=f(v),\qquad\forall~{}f\in\mathcal{H}_{K}.$ (11)
Equipped with $\mathcal{H}_{K}$, we adopt KRR by regressing the response $\mathbf{Y}=(Y_{1},\ldots,Y_{n})^{\top}$ onto the predicted inputs $\widehat{g}(X_{1}),\ldots,\widehat{g}(X_{n})$ by solving the following optimization problem:
@@
Thanks to the reproducing property in (11), the representer theorem (Kimeldorf and Wahba, 1971) states that the minimization in (12) admits a closed-form solution:
$\widehat{f}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\widehat{\alpha}_{i}K_{\widehat{g}(X_{i})}$ (13)
@@
Our theory in Section 3 requires that $\widehat{g}$ be constructed independently of the training data $\mathcal{D}$. This independence simplifies our analysis, especially given the technical complexity of the current framework. Moreover, such independence can lead to better results in many statistical problems, including smaller prediction or estimation errors and weaker conditions – a phenomenon has been observed in various problems, such as estimating the optimal instrument in sparse high-dimensional instrumental variable models (Belloni et al., 2012), inferring a low-dimensional parameter in the presence of high-dimensional nuisance parameters (Chernozhukov et al., 2018), and performing discriminant analysis using a few retained principal components of high-dimensional features (Bing and Wegkamp, 2023). On the other hand, in many applications such as the examples mentioned in the Introduction, auxiliary data without labels or responses are readily available or at least easy to obtain (Bengio et al., 2013). When this is not the case, one can still use a procedure called $k$-fold cross-fitting (Chernozhukov et al., 2018). The 2-fold version of this method first splits the data into equal parts, computes $\widehat{g}$ on one subset, and then performs the regression step using the other subset, before switching their roles. The final prediction is then made by the average of the two resulting predictors.
3 Theoretical guarantees of KRR with predicted feature inputs
In this section we present our theoretical guarantees of the KRR estimator $\widehat{f}$ in (12) that is based on a generic predictor $\widehat{g}:\mathcal{X}\to\mathcal{Z}$, constructed independently of $\mathcal{D}$. The quantity of our interest is the excess risk of the prediction function $\widehat{f}\circ\widehat{g}:~{}\mathcal{X}\to\mathbb{R}$, given by
$\mathcal{E}\bigl{(}\widehat{f}\circ\widehat{g}\bigr{)}:=\mathbb{E}\bigl{[}Y-(\widehat{f}\circ\widehat{g})(X)\bigr{]}^{2}-\sigma^{2}.$ (16)
Henceforth, the expectation $\mathbb{E}$ is taken with respect to a new pair $(Y,X)$ that is independent of, and generated from the same model as, the training data $\mathcal{D}$. Since $\mathcal{E}(\widehat{f}\circ\widehat{g})$ is a random quantity depending on $\mathcal{D}$ and $\widehat{g}$, our goal is to establish its rate of convergence in probability. We start by decomposing the excess risk into three interpretable terms.
@@ 3.1 Decomposition of the excess risk
For any deterministic $f\in\mathcal{L}^{2}(\rho)$, we proved in Section B.1 that
@@
We do not assume $f^{*}\in\mathcal{H}_{K}$ but rather the weaker 1. If the former holds, we immediately have $f_{\mathcal{H}}=f^{*}$. Since $\mathcal{H}_{K}$ is convex, existence of $f_{\mathcal{H}}$ also ensures its uniqueness (Cucker and Smale, 2002). As discussed in Rudi and Rosasco (2017), existence of $f_{\mathcal{H}}$ can be ensured if $\mathcal{H}_{K}$ in 1 is replaced by $\{f\in\mathcal{H}_{K}:\|f\|_{K}\leq R\}$ for any finite radius $R>0$. Alternatively, we can choose $f_{\mathcal{H},\lambda}=\operatorname*{arg\,min}_{f\in\mathcal{H}_{K}}\{\|f-f^{*}\|_{\rho}^{2}+\lambda\|f\|_{K}^{2}\}$ which, for $\lambda>0$, is always unique and exists (Cucker and Smale, 2002). Notably, our results presented below can be applied to any function $f\in\mathcal{H}_{K}$ in replace of $f_{\mathcal{H}}$, including these mentioned examples.
By choosing $f=f_{\mathcal{H}}$ in (17), the second term on the right-hand side (RHS) of (17) represents the irreducible error from two different sources. To see this, by introducing
$\Delta_{\widehat{g}}:=\|K_{Z}-K_{\widehat{g}(X)}\|_{K}^{2},$ (18)
we have that for any $\theta\geq 1$, the elementary inequality $(a+b)^{2}\leq xa^{2}+b^{2}/x$ for $x>0$,
$~{}=~{}(1+\theta)~{}\mathbb{E}\left\langle f_{\mathcal{H}},K_{Z}-K_{\widehat{g}(X)}\right\rangle_{K}^{2}+{1+\theta\over\theta}\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ by (11)
$~{}\leq~{}(1+\theta)\|f_{\mathcal{H}}\|_{K}^{2}~{}\mathbb{E}\Delta_{\widehat{g}}+{1+\theta\over\theta}\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}.$ (19)
The last step uses the Cauchy-Schwarz inequality. The quantity $\mathbb{E}\Delta_{\widehat{g}}$ is referred to as the kernel-related latent error, representing the error in using $\widehat{g}(X)$ to predict the latent factor $Z$, composited with the kernel function $K$. The term $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ on the other hand represents the approximation error due to model misspecification when $f^{*}\notin\mathcal{H}_{K}$. We provide detailed discussion on both terms after stating our main result in Theorem 2 below. Going back to the risk decomposition in (17), the remaining term with $f_{\mathcal{H}}$ in place of $f$, $\mathbb{E}[(Y-(\widehat{f}\circ\widehat{g})(X))^{2}-(Y-(f_{\mathcal{H}}\circ\widehat{g})(X))^{2}]$, represents the prediction risk of $\widehat{f}\circ\widehat{g}$ relative to $f_{\mathcal{H}}\circ\widehat{g}$. Analyzing this term is the main challenge in our analysis, and is detailed in the next section.

### v2

@@ 1.2 Our contributions
A new framework for analyzing KRR with predicted feature inputs.
Along with our new risk bounds comes a novel approach to analyzing the KRR with predicted inputs. The framework we develop offers two key contributions: (1) it improves existing analyses in the classical setting by accommodating arbitrary model misspecification and allowing both the response variable and the RKHS to be unbounded; (2) it is essential for handling predicted feature inputs, enabling us to derive risk bounds without imposing any assumptions on the proximity of $\widehat{g}(X)$ to $Z$.
To accomplish the first point, we develop a new approach that combines the integral operator method with the local Rademacher complexity based empirical process theory from Bartlett et al. (2005), along with a careful reduction argument. This hybrid analysis is necessary because the integral operator approach cannot easily handle model misspecification, while the empirical process theory requires the boundedness of both the response variable and the RKHS. Specifically, as seen in the decomposition (9), the key is to bound from above $\mathbb{E}[\ell_{\widehat{f}\circ\widehat{g}}(Y,X)]$. Since $\sup_{f\in\mathcal{H}_{K}}\|f\|_{K}$ could be unbounded, we first apply a reduction argument, showing that bounding $\mathbb{E}[\ell_{\widehat{f}\circ\widehat{g}}(Y,X)]$ can be reduced to controlling the empirical process
@@
where $\uline{\mathcal{F}_{b}}$ is a bounded subset of $\mathcal{H}_{K}$, given by Eq. 32 of Section 3.4. We use $\mathbb{E}_{n}$ in this paper to denote the expectation with respect to the empirical measure of $n$ i.i.d. samples. Since $Y$ could also be unbounded, we need to bound the cross-term $\mathbb{E}_{n}[\epsilon(f\circ\widehat{g})(X)-\epsilon(f_{\mathcal{H}}\circ\widehat{g})(X)]$ uniformly over $f\in\uline{\mathcal{F}_{b}}$. This is accomplished by employing both the empirical process theory based on local Rademacher complexity from Bartlett et al. (2005) and the integral operator approach, as detailed in Section A.2.3; see also Lemma 7. Our analysis only requires moment conditions on the regression error $\epsilon$. Finally, we again apply the local Rademacher complexity technique of Bartlett et al. (2005) to control the remaining bounded components of the empirical process (11). Notably, our hybrid approach extends to learning algorithms in (1) that use convex losses beyond the squared loss, as illustrated in Section 5.
@@
Another fundamental difficulty in our analysis arises from handling predicted feature inputs. As noted in the end of Section 1.1, the primary challenge lies in establishing a precise connection between the local Rademacher complexity and the kernel complexity function $R(\delta)$ in the presence of errors in predicting the features. The existing analysis in Mendelson (2002) only establishes a connection to $R_{x}(\delta)$, which depends on the probability measure $\rho_{x}$ of $\widehat{g}(X)$. And it is challenging to relate $R_{x}(\delta)$ with $R(\delta)$ without imposing restrictive assumptions on the closeness of $\rho_{x}$ and $\rho$. To circumvent this issue, rather than working with the population-level local Rademacher complexity and $R_{x}(\delta)$—both of which rely on $\rho_{x}$—we instead focus on their empirical counterparts. The first step is to relate the population-level local Rademacher complexity to its empirical counterpart where we borrow existing concentration inequalities in Boucheron et al. (2000, 2003); Bartlett et al. (2005). By adapting the argument in Bartlett et al. (2005, Lemma 6.6), we further link the empirical local Rademacher complexity to $\widehat{R}_{x}(\delta)$, the empirical counterpart of $R_{x}(\delta)$. The next step is to connect $\widehat{R}_{x}(\delta)$ to $\widehat{R}(\delta)$, with the latter being the empirical counterpart of $R(\delta)$. This is the most challenging step, as $\widehat{R}_{x}(\delta)$ and $\widehat{R}(\delta)$ depend, respectively, on the empirical integral operators associated with $\widehat{g}(X_{i})$ and $Z_{i}$ for $i\in[n]$. Our proof reveals that this requires to control the spectral difference between two $n\times n$ kernel matrices with entries equal to $K(\widehat{g}(X_{i}),\widehat{g}(X_{j}))$ and $K(Z_{i},Z_{j})$ for $i,j\in[n]$. As accomplished in Lemmas 11 and 12 of Section A.3.3, this is highly non-trivial as we do not impose any requirements on the proximity of $\widehat{g}(X_{i})$ to $Z_{i}$. Finally, we establish sharp concentration inequalities between $\widehat{R}(\delta)$ and its population-level counterpart, $R(\delta)$, to close the loop of relating the local Rademacher complexity to $R(\delta)$. Completing the loop requires a series of technical lemmas, collected in Lemmas 9, 10, 11, 12 and 13 of Sections A.3.1, A.3.2, A.3.3 and A.3.4. Developing this proof strategy is part of our contribution, and we provide a more detailed discussion in Section 3.4.3. For illustration, Figure 1 below depicts our roadmap of relating the local Rademacher complexity to $R(\delta)$, represented by the solid arrow.
@@
Application to factor-based non-parametric regression models.
Our third contribution is to provide a complete analysis of using KRR with feature inputs predicted by PCA under the factor-based nonparametric regression models (3) and (4). As an application of our general theory, together with the existing results on high-dimensional factor models, we state in Corollary 5 of Section 4.2 an explicit upper bound of the excess risk of the KRR with inputs predicted by PCA. In Theorem 2 of Section 4.2, we further establish a matching minimax lower bound of the excess risk under models (3) and (4), thereby demonstrating the minimax optimality of KRR using the leading PCs.
@@
This paper is organized as follows. The algorithm of KRR with predicted inputs is stated in Section 2. Theoretical results of analyzing KRR with predicted inputs are stated in Section 3. The risk decomposition is given in Section 3.1 while the risk bounds along with the main assumptions are stated in Section 3.2. Risk bounds for specific kernels are given in Section 3.3 while in Section 3.4 we explain the proof sketch and highlight the main technical difficulties. In Section 4 we apply our general result to factor-based nonparametric regression models. In Section 5 we extend our analysis and derive risk bounds for general convex loss functions. Simulation studies and real data analysis that corroborate our theoretical findings are stated in Sections 6 and 7, respectively. All proofs are deferred to the Appendix.
Notation. For any $a,b\in\mathbb{R}$, we write $a\wedge b=\min\{a,b\}$ and $a\vee b=\max\{a,b\}$. For any integer $d$, we let $d]=\{1,\ldots,d\}$. For any symmetric, semi-positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we use $\lambda_{1}(Q)\geq\lambda_{1}(Q)\geq\ldots\geq\lambda_{d}(Q)$ to denote its eigenvalues. For any matrix $A$, we use $\|A\|_{\rm op}$ and $\|A\|_{\text{F}}$ to denote its operator norm and Frobenius norm, respectively. We use $\mathbf{I}_{d}$ to denote the $d\times d$ identity matrix. For $d_{1}\geq d_{2}$, we write $\mathbb{O}_{d_{1}\times d_{2}}$ for matrices with orthonormal columns. The inner product and the endowed norm in the Euclidean space are denoted as $\langle\cdot,\cdot\rangle_{2}$ and $\|\cdot\|_{2}$, respectively. For any function $f:\mathcal{Z}\to\mathbb{R}$, $\|f\|_{\infty}=\sup_{z\in\mathcal{Z}}|f(z)|$. For any two sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C$ such that $a_{n}\leq Cb_{n}$ for all $n$. We write $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. In this paper, we use $c,c^{\prime},C$ and $C^{\prime}$ to denote constants whose values may vary line by line unless otherwise stated.
@@ 2 Kernel ridge regression with predicted feature inputs
In this section, we state the algorithm for Kernel Ridge Regression (KRR) with predicted feature inputs. Recall that $\mathcal{D}=\{(Y_{i},X_{i})\}_{i=1}^{n}\in(\mathbb{R},\mathcal{X})^{n}$ contains the training data. Let $\widehat{g}:\mathcal{X}\to\mathcal{Z}$ be any measurable function used to predict the features $Z_{i}\in\mathcal{Z}$. For instance, $\widehat{g}$ could be obtained from one of the feature extraction or representation learning approaches mentioned in the Introduction. The predicted feature inputs are $\widehat{g}(X_{1}),\ldots,\widehat{g}(X_{n})$.
Let $K:{\mathcal{Z}\times\mathcal{Z}}\to\mathbb{R}$ be some pre-specified kernel function and denote by $\mathcal{H}_{K}$ the Reproducing Kernel Hilbert Space (RKHS) induced by $K$. Write the equipped inner product in $\mathcal{H}_{K}$ as $\langle\cdot,\cdot\rangle_{K}$ and the endowed norm as $\|\cdot\|_{K}$. We propose to regress the response $\mathbf{Y}=(Y_{1},\ldots,Y_{n})^{\top}$ onto $\widehat{g}(X_{1}),\ldots,\widehat{g}(X_{n})$ by solving the following optimization problem:
@@
By letting $K_{v}:=K(v,\cdot)\in\mathcal{H}_{K}$ for any $v\in\mathcal{Z}$, the reproducing property asserts that $\langle f,K_{v}\rangle_{K}=f(v)$ for all $f\in\mathcal{H}_{K}$. As a result, the representer theorem (Kimeldorf and Wahba, 1971) states that the minimization in (12) admits a closed-form solution:
$\widehat{f}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\widehat{\alpha}_{i}K_{\widehat{g}(X_{i})},$ (13)
@@
Our theory in Section 3 requires that $\widehat{g}$ be constructed independently of the training data $\mathcal{D}$. This independence simplifies our analysis, especially given the technical complexity of the current framework. Moreover, such independence can lead to better results in many statistical problems, including smaller prediction or estimation errors and weaker conditions – a phenomenon has been observed in various problems, such as estimating the optimal instrument in sparse high-dimensional instrumental variable models (Belloni et al., 2012), inferring a low-dimensional parameter in the presence of high-dimensional nuisance parameters (Chernozhukov et al., 2018), and performing discriminant analysis using a few retained principal components of high-dimensional features (Bing and Wegkamp, 2023). See, also, our simulation studies in Section 6. On the other hand, in many applications such as the examples mentioned in the Introduction, auxiliary data without labels or responses are readily available or at least easy to obtain. When this is not the case, one can still use a procedure called $k$-fold cross-fitting (Chernozhukov et al., 2018). The 2-fold version of this method first splits the data into equal parts, computes $\widehat{g}$ on one subset, and then performs the regression step using the other subset, before switching their roles. The final prediction is then made by the average of the two resulting predictors.
3 Theory of KRR with predicted feature inputs
In this section we present our theoretical guarantees of the KRR estimator $\widehat{f}$ in (12) that is based on a generic predictor $\widehat{g}:\mathcal{X}\to\mathcal{Z}$, constructed independently of $\mathcal{D}$. The quantity of our interest is the excess risk of the prediction function $\widehat{f}\circ\widehat{g}:~\mathcal{X}\to\mathbb{R}$, given by (5). Since the excess risk is random depending on $\mathcal{D}$ and $\widehat{g}$, our goal is to establish its rate of convergence in probability. We start by decomposing it into three interpretable terms.
@@ 3.1 Decomposition of the excess risk
For any deterministic $f\in\mathcal{L}^{2}(\rho)$, we prove in Section A.1 that
@@
1 is weaker than assuming $f^{*}\in\mathcal{H}_{K}$. If the latter holds, we immediately have $f_{\mathcal{H}}=f^{*}$. Since $\mathcal{H}_{K}$ is convex, existence of $f_{\mathcal{H}}$ also ensures its uniqueness (Cucker and Smale, 2002). As discussed in Rudi and Rosasco (2017), existence of $f_{\mathcal{H}}$ can be ensured if $\mathcal{H}_{K}$ in 1 is replaced by $\{f\in\mathcal{H}_{K}:\|f\|_{K}\leq R\}$ for any finite radius $R$. Alternatively, we can choose $f_{\mathcal{H},\lambda}=\operatorname*{arg\,min}_{f\in\mathcal{H}_{K}}\{\|f-f^{*}\|_{\rho}^{2}+\lambda\|f\|_{K}^{2}\}$ which, for $\lambda>0$, is always unique and exists (Cucker and Smale, 2002). Notably, our results apply to any function $f\in\mathcal{H}_{K}$, in replace of $f_{\mathcal{H}}$, including these examples.
By choosing $f=f_{\mathcal{H}}$ in (16), the second term on the right-hand side of (16) represents the irreducible error from two sources. Indeed, we prove in Section A.1 that for any $\theta\geq 1$,
$\mathbb{E}\left[f^{*}(Z)-(f_{\mathcal{H}}\circ\widehat{g})(X)\right]^{2}$ $~\leq~(1+\theta)\|f_{\mathcal{H}}\|_{K}^{2}~\mathbb{E}\Delta_{\widehat{g}}+{1+\theta\over\theta}\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}.$ (17)
The quantity
$\mathbb{E}\Delta_{\widehat{g}}:=\mathbb{E}\|K_{Z}-K_{\widehat{g}(X)}\|_{K}^{2}$ (18)
is referred to as the kernel-related latent error, representing the error in using $\widehat{g}(X)$ to predict the latent factor $Z$, composited with the kernel function $K$. The term $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ represents the approximation error due to model misspecification when $f^{*}\notin\mathcal{H}_{K}$. We provide detailed discussion on both terms after stating our main result in Theorem 1 below. Back to the decomposition in (16), the first term of the right hand side, with $f_{\mathcal{H}}$ in place of $f$, represents the prediction risk of $\widehat{f}\circ\widehat{g}$ relative to $f_{\mathcal{H}}\circ\widehat{g}$. Its analysis is the main challenge and is detailed in the next section.